Prove by mathematical induction

2n>(n+1)! For all integars n>=2


1
Expert's answer
2021-02-24T07:47:17-0500

The correct inequality is "2^n<(n+1)!"

Basis of induction: n=2.

"2^n = 4 < (2+1)! = 6" is true.

Assume now that the inequality is proved for n=N-1. Then it is correct for n=N:

"(N+1)! = (N+1)\cdot N!>(N+1)\cdot 2^{N-1} = 2\cdot 2^{N-1}=2^N"

Therefore, the statement is proved by the principle of the mathematical induction.